OpenStudy (anonymous):

|6v+15|=3 I got the answer, v=-2, is this right??

OpenStudy (anonymous):

For one of them, yes.

OpenStudy (shamil98):

6v + 15 =3 6v + 15 = -3

OpenStudy (anonymous):

Yea, what he said lol

OpenStudy (shamil98):

6v =-12 v = -2 6v + 15 = -3 6v = -18 v = -3